OpenStudy (anonymous):

find the derivative of 3 cosx cotx

OpenStudy (anonymous):

I just don't understand why the -sinx was taken out of

OpenStudy (anonymous):

-sinx is the derivative of cosx

OpenStudy (anonymous):

but why was cosx substracted from -sinx

OpenStudy (anonymous):

\[\frac{ d }{ dx }\cot x=-\csc ^{2}x\] \[\frac{ d }{ dx }\cos x=-\sin x\]